OpenStudy (anonymous):

Find the quotient. Simplify if possible. 7over9÷ 5over6 =

OpenStudy (anonymous):

who ever gives me the answer you get a medal :D

OpenStudy (kenljw):

7/9 -5/6 = 1/3 (7/3 + 5/2) = (1/3) (14+15)/6 = 29/18

ganeshie8 (ganeshie8):

its a division symbol !

OpenStudy (kenljw):

14 - 15 therefore -1/18

ganeshie8 (ganeshie8):

\(\large \frac{7}{9} \div \frac{5}{6}\) flip the right ratio, and change it into multiplication : \(\large \frac{7}{9} \times \frac{6}{5}\)

OpenStudy (kenljw):

my error sorry

OpenStudy (anonymous):

A.35over63 B.42over45 C.14over15 D.1and1over14

ganeshie8 (ganeshie8):

\(\large \frac{7}{\cancel{9}^3} \times \frac{\cancel{6}^2}{5}\)

ganeshie8 (ganeshie8):

multiply thru and see what you get

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

thats right ! good work !!

OpenStudy (anonymous):

ok here's you medal :)

ganeshie8 (ganeshie8):

lol ty :)

OpenStudy (anonymous):

I have one more left the last medal of the night cheak it out when I post it

ganeshie8 (ganeshie8):

ha

OpenStudy (kenljw):

a/b/c/d = ad/bc top times bottom divided by product of those between 7/9/5/6 = 7X6/9X5 = 7X2/3X5=14/15